Profit or Loss on New Stock Issue

Security Brokers Inc. specializes in underwriting new issues by small firms. On a recent offering of Beedles Inc., the terms were as follows:

Price to public: $5 per share
Number of shares: 3 million
Proceeds to Beedles: $14,000,000

The out-of-pocket expenses incurred by Security Brokers in the design and distribution of the issue were $490,000. What profit or loss would Security Brokers incur if the issue were sold to the public at the following average price?

  1. $5 per share? Use minus sign to enter loss, if any.
    $
  2. $6.75 per share? Use minus sign to enter loss, if any.
    $
  3. $4.25 per share? Use minus sign to enter loss, if any.
Answer #1

a. Price to Public = $5 per share

Number of Shares = 3 million

Proceeds to Security Brokers = 5*3 = $15million

Expenses = $490000

Net Proceeds = $15million - $490000 = $14.51 million

Proceeds to Beedles = $14 million

Profit = $14.51 million -$14million = $0.51 million

b.

Price to Public = $6.75 per share

Number of Shares = 3 million

Proceeds to Security Brokers = 6.75*3 = $20.25million

Expenses = $490000

Net Proceeds = $20.25million - $490000 = $19.76 million

Proceeds to Beedles = $14 million

Profit = $19.76 million -$14million = $5.76 million

c.

Price to Public = $4.25 per share

Number of Shares = 3 million

Proceeds to Security Brokers = 4.25*3 = $12.75million

Expenses = $490000

Net Proceeds = $12.75million - $0.49 = $12.26 million

Proceeds to Beedles = $14 million

Loss = $14 million -$12.26million = $1.74 million